# First-Order Vortex Lattice Melting in Bilayer Ice: A Monte Carlo Method Study

### v1

@@ IV.4.4 First-order transition between the two vortex lattice phases
Similar to the atomic solid phase in the classical limit [42, 43, 44, 45] for the BH model, the phase transition between 1/2 VL and 2/3 VL phases, should be of the first order. The exact boundary between these two phases can be obtained analytically by comparing the free energy $F=E-TS$ of the both phases.
In tradition, the temperature is set to be $T=1$. By carefully checking, the entropy for the both phase are $S_{1/2}=\ln{64}$ and $S_{2/3}=\ln{36}$, respectively. The entropy does not depend on the lattice size, and therefore, the average entropy for persite denoted as $S/L^{2}/2$ should be zero in the thermal dynamical limit $L\rightarrow\infty$. The energies of the two phases are defined as follows:
@@
where $-L^{2}\ln{a}-L^{2}\ln{b}$ and $-\frac{2}{3}L^{2}\ln{a}-\frac{2}{3}L^{2}\ln{b}-\frac{2}{3}L^{2}\ln{c}$ are the vertex energy for the both phases, and $-\frac{1}{2}L^{2}\ln{V}$ and $-\frac{2}{3}L^{2}\ln{V}$ are the energies for the vortex. Let
@@
with keeping $a=b$, the reduced analytical expression becomes
@@
Numerically simulations also are performed to confirm the first order transition between the vertex phases. In Fig. 13 (a), the distribution of $\rho_{vx}$ for the sizes $L=12,24$ are obtained. In Fig. 13 (b), $\rho_{vx}$ versus $L$ are plotted. The gray data are MC results with 20 independent bins. The theoretical $a_{c}$, are marked by the red line for guide eyes.
@@
In the regime of a strong first-order transition, the cluster algorithm encounters the issue of ergodicity. As a result, we have also incorporated the Metropolis type of short-loop update scheme. In the $z$ direction, the loops connecting two layers can be flipped with a certain probability without violating the “ice” rule for each involved vertex. In order to keep the “two-in two-out” constraint, we flip eight legs as shown in Fig. 13 (c), where the number of 1, 2, 3, 4 mean two of the four legs for each vertex.
In the actual simulation, after each round of loop updates, we will check if there are any short loops present. If a short loop is found, we will attempt to flip it.
Figure 13: Signal of the first-order phase transition. (a) The double-peaked of the distribution of $\rho_{vx}$, with $V=7$, $a=1.68$ (b) $\rho_{vx}$ versus $a$ (c) a short loop in the vertical direction, and the numbers 1, 2, 3, 4 represent the legs. (d) The Metropolis type of short-loop updating.
V Discussion and conclusion
@@
Our study contrasts the square lattice 6V model, where the AFM to disordered phase transition is governed by the BKT mechanism, with the bilayer 6V model, characterized by a conventional Ising phase transition due to Z2 symmetry breaking. We find that the transitions from vortex lattice phases to the disorder phase and between different vortex lattice phases are both of first order. These insights advance our understanding of the phase transitions present in layered water ices, contributing to the broader comprehension of complex systems in physics.
@@
(i) As an initial investigation, we have assumed $a=b$, and the case where $a\neq b$ has not been studied yet. The exploration of vortex glass induced by random values of $a$, $b$, and $V$ worthwhile. Furthermore, the disorder we observed can potentially be further classified into disordered structures with vortices and disordered structures without vortices.
@@
(iii) Although we are not clear about how water ice regulates the ratios of $a$, $b$, and $c$ types of vertexes, it is possible to artificially adjust them in the case of artificial spin ice [3], which has many rich physics.
(iv) Regarding the numerical methods for studying this model, not only conventional MC methods are suitable, but tensor network methods [51] are also well-suited for exploring this model. There are existing literature studies that have employed tensor network methods to investigate similar models such as the dimer model and ice model. However, in our case, we introduce a slightly more complex factor by incorporating vortex weighting, which results in an increase in the bond dimensions of the tensor elements.
Acknowledgements We thank Vladimir Korepin and Tzu-Chieh Wei for their inspiration regarding the six-vertex model, and Youjin Deng and Chengxiang Ding for carefully reading and commenting the manuscript. This work was supported by the Hefei National Research Center for Physical Sciences at the Microscale (KF2021002), and Shanxi Province Science Foundation (Grants No: 202303021221029 and No:202103021224051).
Appendix A Configurations with the $a$, $b$, $c$ types vertexes and vortexes
In Figs. A2 (a)-(d), the snapshots of the configurations with vertexes and vortexes are shown. The parameters chosen correspond to the positions marked as ①, ②, ③, ④ in Fig. A2 (e).
In the 2/3 vortex lattice phase, there not only exist vortex orders but also AFM orders with a length of 3 in the $x$-direction Fig. A2 (f). If $y==1$, $S_{y,x}=1$ for $x\in[1,2,3]$ and $S_{y,x}=-1$ for $x\in[1,2,3]$. The definition to quantify such pattern is as follows
$M_{3}=\sum_{i=1}^{L}\left|\sum_{j=1}^{L}(-1)^{\left\lfloor(j-1)/3\right\rfloor}S_{i,j}\right|/N,$ (37)
where $\lfloor(j-1)/3\rfloor$ means replacing 1,2,3,4,5 and 6 to 0,0,0,1,1, and 1. Similarily, for the 1/2 VL lattice, there is FM order in the $x$ direction and the quantitiy is defined in Eq. 10.
@@ Appendix B Description of the bounce probabilities
In Sec. II.2, Table 1 shows the range of bounce probabilities $p_{a}$, $p_{b}$ and $p_{c}$. In this appendix, more details are described. Corresponding to the 6V model, the general expression of the bounce probability of the loop algorithm is
@@
where the denominator is the sum of the weights of the three vertices after the loop passes through two allowable exit legs and one bounce leg (Fig. A1), and the numerator $w_{i}$ is the weights of the vertices encountered during the loop walk. The subscript $i$ denotes types of $a$, $b$, or $c$. To discuss the range of $p_{i}$, it is only necessary to discuss the maximum and minimum values of Eq. 38, where the variable $w_{c}$ is already fixed to 1. Taking $p_{a}\in(0,1/2)$ in the AFM phase as an example, the weights $w_{a}$ and $w_{b}$ are independently adjustable, but they meet the range $10<a+b<$ 1, and therefore the maxvalue of $p_{a}=a/(a+b+1)\rightarrow a/(a+1)\rightarrow 1-1/(a+1)=1/2$.
Figure A2: Snapshots for different phases. (a) $\rho_{v}=1/2$ striped vortex lattice (b) $\rho_{v}=2/3$ vortex lattice (c) disorder phase without vortex (d) disorder phase with vortex (e) the four snapshots are obtained at parameters marked in ①, ②, ③ and ④. (f) The AFM pattern with minmum unit length 3 on the 2/3 VL lattice.

### v2

@@ IV.4.4 First-order transition between the two vortex lattice phases
Figure 13: Signal of the first order phase transition. (a) The double peaked distribution of $\rho_{vx}$, with $V=7$, $a=1.68$ (b) $\rho_{vx}$ versus $a$ (c) a short loop in the vertical direction, and the numbers 1, 2, 3, 4 represent the legs. (d) The Metropolis type short-loop updating.
Similar to the atomic solid phase in the classical limit [43, 44, 45, 46] for the BH model, the phase transition between 1/2 VL and 2/3 VL phases, should be of first order. The exact boundary between these two phases can be obtained analytically by comparing the free energy $F=E-TS$ of both phases.
By convention, the temperature is set to be $T=1$. By careful checking, the entropy for both phases is respectively $S_{1/2}=\ln{64}$ and $S_{2/3}=\ln{36}$. The entropy does not depend on the lattice size, and therefore, the average entropy for per site denoted as $S/L^{2}/2$ should be zero in the thermal dynamical limit $L\rightarrow\infty$. The energies of the two phases are defined as follows:
@@
where $-L^{2}\ln{a}-L^{2}\ln{b}$ and $-\frac{2}{3}L^{2}\ln{a}-\frac{2}{3}L^{2}\ln{b}-\frac{2}{3}L^{2}\ln{c}$ are the vertex energies for both phases, and $-\frac{1}{2}L^{2}\ln{V}$ and $-\frac{2}{3}L^{2}\ln{V}$ are the energies for the vortex. Let
@@
while keeping $a=b$, the reduced analytical expression becomes:
@@
Numerical simulations are also performed to confirm the first order transition between the vertex phases. In Fig. 13 (a), the distributions of $\rho_{vx}$ for the sizes $L=12,24$ are obtained. In Fig. 13 (b), the curves $\rho_{vx}$ versus $L$ are plotted. The gray data are MC results with 20 independent bins. The theoretical $a_{c}$, are marked by the red line to guide the eyes.
@@
In the regime of a strong first order transition, the cluster algorithm encounters the issue of ergodicity. As a result, we have also incorporated the Metropolis type short-loop update scheme. The short loop in the $z$-direction is composed of four vertices, with each vertex contributing two legs that form the rectangles depicted in Figure 11(c). These pairs of legs are (3,4), (2,3), (1,2) and (4,1), respectively. In order to keep the “two-in-two-out” constraint, we flip eight legs simultaneously with a certain probability without violating the “ice” rule for each involved vertex. During the actual simulation, after each round of loop updates, we check if there are any short loops present. If a short loop is found, we attempt to flip it.
IV.4.5 Effect of temperature gradient field
Figure 14: The distribution of vortex density $\rho_{v}(x,y)$ with temperature gradient field (a) along the $x$ direction and (b) along the $y$ direction. The temperature varies from 1 to 0.45 with an interval of 0.05, with $a=b=1.5,c=1,V=2$ at the site (1,1).
On the single-layer honeycomb lattice, vortex lattice has been achieved using the charged colloidal particle-based ice numerically [52]. The Coulomb repulsion between charged particles favors vortices at low temperatures. Experimentally, a magnetic colloidal particle [53] inside lithographically sculptured double well traps, with tunable repulsive interactions of magnetic moments, also favors vortices.
However, raising the temperature  [52] results in the vortex state melting into a disordered state. Additionally, introducing a magnetic field [54] or an electric field in the plane also easily transitions the vortex state into a ferromagnetic state.
Another factor of the vortex depinning is a temperature gradient field. Here we put a temperature field distributed on the lattice, with the temperature decreasing. The pattern of vortex density $\rho_{v}(x,y)$ becomes striped as shown in Fig. 14 (a)-(b). The vortex density $\rho_{v}(x,y)$ at higher temperature region are larger than that in the lower temperature region. Specifically, the coordinate variables $x_{1},x_{2},\cdots,x_{L}$ are mapped to $T_{1},T_{2},\cdots,T_{L}$, and ultimately to $a_{1},a_{2},\cdots,a_{L}$ by $a_{i}=b_{i}=\text{exp}(\epsilon_{i}/T_{i})$. A similar mapping applies to the vortex weight $V_{i}$.
Note that before applying the temperature gradient field, we set the temperature to 1. Then, we calculate the energy of each vertex in the lattice based on parameters $a$, $b$, $c$ and the Boltzmann weight. When introducing the temperature gradient field, it’s crucial to maintain these energies fixed, as vertices at different positions experience different temperatures.
When an external field is applied, vortices either disappear or form stripes. Without an external field, due to translational invariance, the vortex density is uniformly distributed. The different average distributions of vortex density with and without an external field indicate vortex depinning.
V Experiment realization using particle based ice
The colloidal particle-based ice, as discussed in [55], is another system capable of realizing bilayer ice, in addition to water ice, natural spin ice, and artificial spin ice. The theoretical concept behind colloidal particle ice involves employing elongated optical traps to confine colloidal particles. Each optical trap features two wells where colloidal particles can reside. These arrays of traps can organize into different kinds of lattices [55, 52, 56]. Additionally, the presence of topological charges [57], vortex states [58] and skyrmions patterns [59], boundary effects [60], can be simulated by particle-based ice. More complete reviews are present in Ref. [5].
Figure 15: The method of embedding the A-B stacked bilayer honeycomb lattice into the single-layer substrate (a) The single-layer rendition of two honeycomb lattices, with double-sided arrows indicating their connection. (b) Representation of two different double well traps for the two lattices, each accommodating a single colloidal particle. These traps can be symbolized as bonds. (c) Illustration of the “two-in-two-out” rule pattern, where site B connects four A sites, including one A site from a different lattice. (d) Depiction of crossing bonds and their corresponding double well structures.
Experimentally, bilayer water ice, where the ordering is due to the proton positions [22], is probably better captured by particle-based ice [5]. Particle-based systems could include charged or magnetic colloids interlacing in an array that mimics spin ice and more closely resembles the water ice’s proton ordering despite their different natures [54]. The implementation of a two-well trap array embedded in a substrate has been achieved in references [61, 62, 53], on two-dimensional lattices. However, for a bilayer lattice, achieving the required structure has not yet been accomplished.
Here, we propose imprinting both layers onto a single-layer substrate, as schematically illustrated in Fig. 15 (a). The upper and lower honeycomb lattices are depicted in blue and black, respectively. Interlayer connections are denoted by double arrows. Given that this is an A-B stacked bilayer, the sites on the A sublattice and the B sublattice are connected pairwise.
As depicted in Fig. 15 (b), each edge can be represented as a double well holding only one colloid. Some edges corresponding to the two layers intersect, raising the possibility of particle collision at these intersections. To reduce this, the depths of the two wells can be adjusted differently. The black wells are deep, while the blue ones are shallow. Additionally, a small well between the two blue wells helps the passage of small colloids.
In Fig. 15 (c), the arrangement illustrates four non-crossing bonds. Site B (in the lower layer) connects to four A sites, with the blue A site originating from the upper layer. In Fig. 15 (d), the crossing bonds are depicted. The shallow and deep wells meet at the intersection from different directions. Both the big and small colloids can move within their respective wells.
VI Discussion and conclusion
@@
Our study contrasts the square lattice 6V model, where the AFM to the disordered phase transition is governed by the BKT mechanism, with the bilayer 6V model, characterized by a conventional Ising phase transition due to Z2 symmetry breaking. We find that the transitions from vortex lattice phases to the disorder phase and between different vortex lattice phases are both of first order. This insight advances our understanding of the phase transitions present in layered water ices, contributing to the broader comprehension of complex systems in physics.
@@
(i) As an initial investigation, we assumed $a=b$, and the case where $a\neq b$ has not been studied yet. The exploration of vortex glass induced by random values of $a$, $b$, and $V$ worthwhile. Furthermore, the disorder we observed can potentially be further classified into disordered structures with vortices and disordered structures without vortices.
@@
(iii) Although we are not clear about how water ice regulates the ratios of $a$, $b$, and $c$ types of vertices, it is possible to artificially adjust them in the case of artificial spin ice [3], which is rich in physical properties.
(iv) Regarding the numerical methods for studying this model, not only are conventional MC methods suitable, but tensor network methods [63] are also well-suited for exploring this model. There are existing literature studies that have employed tensor network methods to investigate similar models such as the dimer model and the ice model. However, in our case, we introduce a slightly more complex factor by incorporating vortex weighting, which results in an increase in the bond dimensions of the tensor elements.
Acknowledgments We thank Vladimir Korepin and Tzu-Chieh Wei for their inspiration regarding the six-vertex model, and Youjin Deng, Chengxiang Ding, and tcr C. Scott of RWTH-Aachen University for commenting on the manuscript, and valuable suggestions from the reviewers. This work was supported by the Hefei National Research Center for Physical Sciences at the Microscale (KF2021002), and the Shanxi Province Science Foundation (Grants No: 202303021221029 and No:202103021224051).
Appendix A Configurations with the $a$, $b$, $c$ types vertices and vortices
In Figs. A2 (a)-(d), the snapshots of the configurations with vertices and vortices are shown. The parameters chosen correspond to the positions marked as ①, ②, ③, ④ in Fig. A2 (e).
In the 2/3 vortex lattice phase, not only do vortex orders exist but also AFM orders with a length of 3 in the $x$-direction Fig. A2 (f). If $y==1$, $S_{y,x}=1$ for $x\in[1,2,3]$ and $S_{y,x}=-1$ for $x\in[1,2,3]$. The definition to quantify such a pattern is as follows:
$M_{3}=\sum_{i=1}^{L}\left|\sum_{j=1}^{L}(-1)^{\left\lfloor(j-1)/3\right\rfloor}S_{i,j}^{x}\right|/N,$ (37)
where $\lfloor(j-1)/3\rfloor$ means replacing 1,2,3,4,5 and 6 to 0,0,0,1,1, and 1. Similarly, for the 1/2 VL lattice, there is a FM order in the $x$ direction and the quantity is defined in Eq. 10.
@@ Appendix B Description of the bounce probabilities
In Sec. II.2, Table 1 shows the range of bounce probabilities $p_{a}$, $p_{b}$ and $p_{c}$. In this appendix, more details are described. Corresponding to the 6V model, the general expression of the bounce probability of the loop algorithm is:
@@
where the denominator is the sum of the weights of the three vertices after the loop passes through two allowable exit legs and one bounce leg (Fig. A1), and the numerator $w_{i}$ are the weights of the vertices encountered during the loop walk. The subscript $i$ denotes types of $a$, $b$, or $c$. To discuss the range of $p_{i}$, it is only necessary to discuss the maximum and minimum values of Eq. 38, where the variable $w_{c}$ is already fixed to 1. Taking $p_{a}\in(0,1/2)$ in the AFM phase as an example, the weights $w_{a}$ and $w_{b}$ are independently adjustable, but they meet the range $10<a+b<$ 1, and therefore the maximum value of $p_{a}=a/(a+b+1)\rightarrow a/(a+1)\rightarrow 1-1/(a+1)=1/2$.
Figure A2: Snapshots for different phases. (a) $\rho_{v}=1/2$ striped vortex lattice (b) $\rho_{v}=2/3$ vortex lattice (c) disordered phase without vortex (d) disordered phase with vortex (e) the four snapshots are obtained at parameters marked in ①, ②, ③ and ④. (f) The AFM pattern with minimum unit length 3 on the 2/3 VL lattice.